# Large Language Models for Crash Detection in Video: A Survey of Methods, Datasets, and Challenges

### v1

@@ VII-A Data Scarcity
LLM and VLM models require large, diverse datasets of labeled video-caption pairs to achieve robust performance in crash detection. However, such datasets are scarce, particularly for crash-specific scenarios, due to the rarity of crash events and the complexity of annotating video data [38]. While recent large-scale datasets like SHRP 2 NDS [42] have significantly improved data availability, they still represent a fraction of the data needed to cover all possible real-world scenarios. Existing datasets, such as the Dashcam Accident Dataset (DAD) and UCF-Crime, contain limited crash-related videos, often with coarse or incomplete annotations [38, 41]. For instance, DAD includes only 1,500 annotated videos, insufficient for training large-scale multimodal models like Video-LLaMA [27]. Synthetic datasets, such as those used in CrashLLM, attempt to address this by generating video-caption pairs, but they may lack the variability of real-world traffic conditions [33]. Data scarcity also exacerbates dataset imbalance, where non-crash videos dominate, leading to biased model predictions [11]. Efforts to create comprehensive, crash-specific datasets are critical to improving model generalization and performance [39].
@@ VII-B Multimodal Alignment
Achieving temporal synchronization between video frames and language descriptions is a significant challenge, particularly in crash detection scenarios involving occlusions, low framerates, or complex traffic dynamics [29]. Multimodal alignment requires models to correlate visual events (e.g., a collision) with corresponding textual descriptions (e.g., “A car collided with a truck”), ensuring temporal and semantic consistency [27]. For example, frameworks like VERA struggle to align frame-level visual features with textual captions under occlusions, where key crash indicators may be obscured [29]. Low framerates, common in traffic surveillance videos, further complicate alignment by reducing temporal resolution, as noted in studies of classical optical flow methods [4]. Cross-attention mechanisms, as used in VAD-LLaMA, mitigate this by modeling long-range dependencies, but they increase computational complexity [34]. Developing efficient alignment techniques that handle occlusions and variable framerates remains an open research problem [11].
@@ VII-C Reasoning and Explainability
While LLMs generate fluent crash reports, their outputs may be factually incorrect or inconsistent with visual inputs, a phenomenon known as hallucination [30]. This poses a critical challenge for crash detection, where accurate and interpretable outputs are essential for applications like autonomous driving and post-event analysis [33]. For instance, CrashLLM’s generative descriptions may inaccurately attribute crash causes (e.g., “sudden braking” instead of “lane change”) if visual cues are ambiguous [33]. Ensuring consistency between LLM-generated reports and visual evidence requires robust grounding techniques, such as those explored in VERA, which uses LLaVA-1.5 to verbalize crash rationales [29]. However, even advanced models struggle with causal reasoning under complex traffic scenarios, as noted in studies of multimodal forecasting [35]. Improving explainability through techniques like scene-graph grounding or causal intervention modules is a key open problem to enhance trust in LLM-based systems [26, 11].
@@ VII-D Real-time Constraints
Real-time crash detection is critical for applications like accident response and autonomous vehicle decision-making, but LLM-based pipelines often exceed the computational limits required for real-time inference [37]. Large models, such as GPT-4 or Video-LLaMA, demand significant resources, with inference times ranging from hundreds of milliseconds to seconds per video clip, far exceeding the 30–100 ms latency required for real-time systems [30, 27]. For example, while the specific hardware is not detailed, the VADSK paper notes that the foundational models used within Holmes-VAD have inference speeds ranging from approximately 2 to 6 seconds per frame, clearly outside the real-time window for immediate response [58]. This highlights a significant performance gap that needs to be addressed for practical deployment.
Optimizing model architectures (e.g., using lightweight adapters like LoRA [32]), developing efficient temporal modules, and leveraging quantization techniques are essential to meet real-time constraints while maintaining performance [11]. For example, Holmes-VAD’s multi-modal LLM achieves high accuracy but incurs high computational costs, limiting its use in resource-constrained environments [56]. Training-free approaches like LAVAD reduce latency by eliminating fine-tuning, achieving 85.0% AUC with lower overhead, but they may sacrifice precision in complex scenarios [37]. Classical methods, such as frame differencing, offer lower latency but lack the robustness of LLMs [3]. A quantitative comparison of inference times and hardware requirements across models remains a challenge due to a lack of standardized reporting in the literature, but the available data suggests a clear trade-off between model complexity, accuracy, and processing speed.
@@ VII-E Ethical and Fairness Considerations
Beyond technical hurdles, the deployment of LLM-based crash detection systems raises significant ethical and fairness concerns. Datasets such as DAD and UCF-Crime, while foundational, may contain inherent biases by over-representing specific geographic regions, weather conditions, or vehicle types. Models trained on such data may perform poorly in underrepresented scenarios, leading to inequitable safety outcomes. Furthermore, the use of dashcam and surveillance footage introduces profound privacy risks, as this data can capture identifiable information like faces and license plates. The societal impact of large-scale video data collection and analysis requires careful consideration of data anonymization techniques and transparent policies regarding data usage and storage.
Algorithmic fairness is another critical issue. Biased data can lead to biased models, where false positive or negative rates are disproportionately higher for certain demographic groups or environments. While methods like Holmes-VAD aim to address unbiased detection, the risk remains that automated systems could perpetuate or even amplify existing societal biases, with serious consequences for insurance liability and legal accountability. Therefore, developing fair, transparent, and privacy-preserving systems is not just a technical challenge but an ethical imperative for the responsible advancement of this technology.
VII-F Robustness and Generalization Analysis
A critical challenge for LLM-based crash detection models is their ability to generalize to scenarios not well represented in the training data. While models may perform well on benchmarks, their robustness in unseen or out-of-distribution (OOD) scenarios—such as rural roads, adverse weather (e.g., snow, heavy rain), or different geographic regions—remains a significant concern. This limitation poses a risk of overfitting to the visual and contextual cues prevalent in the training data, leading to a sharp degradation in performance when deployed in the real world [59].
VII-F1 Domain Generalization
To address this, domain generalization techniques are essential. These methods aim to learn domain-invariant features that are robust to shifts in data distribution. Techniques such as domain-invariant feature learning, meta-learning, and adversarial domain adaptation can help models generalize from multiple source domains to unseen target domains [60]. Data augmentation, particularly using high-fidelity simulators like CARLA, allows for the creation of diverse training data by simulating varied weather and lighting conditions, which can significantly improve model robustness [61].
VII-F2 Out-of-Distribution (OOD) Detection
Complementary to domain generalization is Out-of-Distribution (OOD) detection, which focuses on identifying whether a given input is from a novel, unseen distribution. In crash detection, this is crucial for flagging rare events (e.g., unusual multi-vehicle pile-ups, crashes involving non-standard vehicles) that the model was not trained to recognize. By equipping models like LAVAD or VERA with an OOD detection mechanism, the system could defer to a human operator or a fail-safe mode when it encounters a scenario it cannot confidently analyze, thereby increasing safety. Methods based on uncertainty estimation are particularly promising for this task, as they can quantify the model’s confidence in its predictions [62].
VII-F3 Cross-Dataset Evaluation and Quantitative Analysis
A rigorous evaluation methodology is necessary to properly assess generalization. This should involve comprehensive cross-dataset evaluation, where models trained on one dataset (e.g., DAD) are tested on another with different characteristics (e.g., UCF-Crime or the more diverse SHRP 2 NDS) [40, 42]. A quantitative analysis of this performance drop, for example, by reporting the decrease in AUC or F1-score, provides a clear measure of a model’s generalization capability. Future work should standardize this practice and ideally present results in a comparative table, analyzing performance drops with statistical significance tests to build a clearer picture of which models are most robust.
VII-F4 Augmentation with Synthetic Data
@@
The challenges of data scarcity, multimodal alignment, reasoning and explainability, real-time constraints, and generalization underscore the complexities of deploying LLM and VLM-based crash detection systems. Data scarcity limits model generalization, necessitating larger, more diverse datasets [39]. Multimodal alignment issues highlight the need for advanced fusion strategies, such as cross-attention, to handle occlusions and low framerates [34]. Reasoning and explainability challenges demand robust grounding techniques to ensure factual accuracy [29]. Real-time constraints require optimized architectures to balance latency and performance [37]. Finally, ensuring robust generalization requires a combination of advanced techniques like domain generalization, OOD detection, and the use of synthetic data, all validated through rigorous cross-dataset evaluation. Addressing these challenges will enhance the reliability and applicability of LLM-based crash detection, building on the advancements discussed in Section VI and paving the way for future research in intelligent transportation systems.
VIII Ethical, Legal, and Societal Implications
Beyond the technical challenges, the deployment of LLM-based crash detection systems carries profound ethical, legal, and societal implications that demand careful consideration. As these technologies become more integrated into safety-critical domains, a proactive approach to addressing their broader impact is essential for responsible innovation and public trust.
VIII-A Bias and Fairness
A primary concern is the risk of algorithmic bias. Datasets used to train crash detection models, such as DAD and UCF-Crime, often over-represent urban environments and specific weather conditions, potentially leading to biased models that underperform in rural or atypical scenarios [63]. This can result in inequitable safety outcomes, where the system is less reliable for certain communities. To mitigate this, researchers propose debiasing techniques like reweighting underrepresented data samples and adversarial training to create more robust and equitable models. Ensuring fairness requires a concerted effort to build diverse and representative datasets that reflect the full spectrum of real-world driving conditions.
VIII-B Privacy Risks
The use of dashcam and surveillance footage for crash detection raises significant privacy concerns, as this data can capture personally identifiable information such as faces and license plates. The large-scale collection and analysis of such video data require robust privacy-preserving measures. Techniques like homomorphic encryption, which allows computation on encrypted data, and differential privacy, which adds statistical noise to data to protect individual identities, are promising solutions [64]. Furthermore, strict adherence to data protection regulations like GDPR and CCPA is necessary to ensure that data is handled responsibly and transparently.
VIII-C Legal Accountability
The increasing autonomy of crash detection systems introduces complex questions of legal accountability. When an LLM-based system like CrashLLM or VERA generates a crash report that is used in insurance claims or legal proceedings, determining liability in the event of an error becomes challenging. Is the responsibility with the vehicle owner, the software developer, or the vehicle manufacturer? A clear legal framework is needed to address these new liability scenarios. Recent surveys on AI safety assurance emphasize the need for new standards and regulations that cover the entire lifecycle of AI systems, from design and testing to deployment and maintenance, to establish clear lines of accountability [65].
VIII-D Societal Impact and Public Trust
The societal impact of LLM-based crash detection includes significant benefits, such as the potential for faster emergency response and a reduction in road fatalities. However, there are also risks, including over-reliance on automated systems and public apprehension. Building public trust is paramount for the successful adoption of these technologies. Surveys on public trust in AI reveal that while citizens see the potential benefits, they also harbor significant concerns about job loss, privacy, and misuse of the technology by authorities [66]. Fostering trust requires transparent communication about how the technology works, its limitations, and the safeguards in place. Engaging the public in the development and oversight process and ensuring that the systems are explainable and accountable will be key to building public confidence [63].
VIII-E Case Study: A Hypothetical Scenario
Consider a hypothetical scenario where an LLM-based crash detection system is widely deployed in a country. The system works exceptionally well in urban areas but has not been sufficiently trained on data from rural regions with unpaved roads and different types of agricultural vehicles. An accident occurs on a rural road involving a tractor, but the system fails to detect it in a timely manner, delaying emergency services. The legal investigation reveals the system’s biased performance. This scenario highlights the severe consequences of dataset bias, leading to legal challenges against the manufacturer for negligence, regulatory fines for deploying an unsafe product, and a significant erosion of public trust in AI-powered safety systems. It underscores the critical need for comprehensive testing, fair representation in training data, and robust regulatory oversight before widespread deployment.
IX Integration with Autonomous Driving Ecosystems
@@
IX-A AV Pipeline Integration
A standard AV software architecture follows a perception-planning-action pipeline [67]. LLM-based crash detection serves as an advanced component of the perception stack. The outputs from a model like VERA or CrashLLM—such as a textual description (”imminent rear-end collision due to sudden braking”) or a classification of risk—can be fed directly into the AV’s decision-making or planning module. This allows the AV to move beyond simple object detection and react to a nuanced understanding of the scene’s dynamics. For example, a high-risk prediction could trigger an emergency braking maneuver or a swerving action, demonstrating a direct link between semantic crash understanding and vehicle control.
IX-B Multi-Sensor Fusion
While cameras are the primary input for the VLM-based methods discussed, their fusion with other sensor modalities like LiDAR and radar is crucial for building a robust perception system for AVs. LiDAR provides precise 3D geometric data, which is less affected by adverse weather, while radar offers reliable velocity information. A multi-sensor fusion approach would combine the semantic, contextual understanding from an LLM/VLM with the geometric and velocity data from LiDAR/radar [68]. For example, a VLM might identify a ”pedestrian looking at their phone,” and this semantic flag could be fused with LiDAR data confirming their trajectory towards the road, leading to a more reliable and timely pre-emptive braking action than either sensor could achieve alone. Extending frameworks like Video-LLaMA to incorporate these additional data streams is a promising research direction.
IX-C Real-time Performance and Latency
@@
IX-D Crash Anticipation as a Proactive Safety Measure
@@
X Quantitative Benchmarking and Ablation Studies
To provide a clearer, more rigorous comparison of the methods discussed, this section proposes a framework for unified quantitative benchmarking and outlines the need for systematic ablation studies.
X-A Cross-Dataset Generalization
A crucial test of robustness is cross-dataset generalization. Models should be trained on one dataset (e.g., CADP) and evaluated on another (e.g., BDD100K) without fine-tuning. This tests the model’s ability to generalize to new environments, camera types, and traffic patterns. Reporting the performance drop in such scenarios would provide a quantitative measure of a model’s generalization capabilities, which is currently a major gap in the literature.
X-B Ablation Studies
To understand the contribution of individual architectural components, systematic ablation studies are needed. For instance, in a model like VERA, one could ablate the language-based reasoning module to quantify its impact on performance compared to using only the visual features. Similarly, for Video-LLaMA, ablating the temporal module would reveal how much the model relies on temporal cues versus single-frame analysis. Such studies are essential for understanding what makes these models effective and for guiding future research [70].
X-C Computational Metrics
For practical deployment, especially in AVs, computational performance is as important as accuracy. Future benchmarking should include a systematic comparison of inference latency (in milliseconds) and computational requirements (e.g., GPU memory, TFLOPs) on standardized hardware. For example, comparing the performance of a heavyweight model like Holmes-VAD against a lightweight, training-free model like LAVAD would highlight the trade-offs between accuracy and efficiency, providing critical insights for real-world applications where resources are constrained [58].
XI Future Directions
The challenges outlined in Section VII and the advancements in LLM-based crash detection methods discussed in Section VI highlight the need for innovative approaches to enhance the robustness, scalability, and applicability of these systems in intelligent transportation systems. This section proposes key future directions to address current limitations and advance the field.
XI-A Synthetic Training Data
The scarcity of labeled video-caption pairs for crash detection, as noted in Section VII, limits the generalization of LLMs and VLMs [38, 39]. To address this, synthetic training data generated through simulation platforms like CARLA or SUMO offers a promising solution [40]. These platforms can create diverse crash scenarios, including rare events (e.g., multi-vehicle collisions under adverse weather), paired with detailed textual narratives (e.g., “A truck skidded on a wet road, colliding with a sedan”). For instance, CARLA’s high-fidelity simulations have been used to generate video data for autonomous driving tasks, which can be extended to crash-specific datasets with annotations for temporal events and causal factors [40]. Synthetic data can augment existing datasets like DAD and UCF-Crime, reducing dataset imbalance and improving model robustness [41]. Recent work on CrashLLM demonstrates the potential of synthetic video-caption pairs, though real-world variability remains a challenge [33]. Future efforts should focus on hybrid datasets combining synthetic and real-world data, leveraging techniques like domain adaptation to ensure transferability [23]. Such approaches can mitigate data scarcity, enabling models to handle diverse traffic conditions and rare crash scenarios [11].
XI-B Video-grounded QA Benchmarks
@@
XI-C Fine-tuned VLMs
@@
While pre-trained VLMs like BLIP-2 and Flamingo offer strong generalization, their performance in crash-specific scenarios is limited by the lack of domain-specific fine-tuning [25, 26]. Fine-tuning VLMs on crash-specific scenes and annotations can enhance their ability to detect and describe complex crash events, addressing multimodal alignment issues [29]. For instance, fine-tuning Video-LLaMA on datasets like DAD or BDD100K, with annotations for crash types, severity, and temporal dynamics, can improve temporal synchronization and feature extraction [27, 40]. Techniques like LoRA (Low-Rank Adaptation) enable efficient fine-tuning of large models, reducing computational overhead while adapting to crash-specific contexts [32]. Recent work on VERA demonstrates the benefits of fine-tuning LLaVA-1.5 for explainable crash detection, achieving high AUC and BLEU scores [29]. Future efforts should focus on curating crash-specific annotations, incorporating temporal models like TimeSformer or SlowFast networks to capture dynamic events [52, 46]. Such fine-tuned VLMs will enhance accuracy and robustness, particularly in challenging scenarios involving occlusions or low framerates, as discussed in Section VII [11].
XI-D Multilingual and Low-resource Models
Most LLM and VLM-based crash detection systems rely on English-language datasets and descriptions, limiting their applicability in non-English-speaking regions [33]. Expanding to multilingual and low-resource models can improve accessibility, enabling crash detection systems to process descriptions in languages like Spanish, Mandarin, or Hindi. Multilingual pre-training, as explored in models like MiniGPT-4, provides a foundation for handling diverse linguistic inputs [45]. For example, fine-tuning Video-LLaMA on multilingual crash datasets, such as those derived from global traffic surveillance, can enable cross-linguistic crash analysis [27]. Low-resource models, optimized for environments with limited computational resources, are equally critical for deployment in developing regions. Techniques like knowledge distillation or training-free approaches, as used in LAVAD, can reduce model complexity while maintaining performance [37]. Initiatives like TrafficLens demonstrate the potential of lightweight models for video-to-text analysis in resource-constrained settings [36]. Future work should prioritize multilingual dataset creation and efficient model architectures, ensuring equitable access to crash detection technologies across diverse linguistic and economic contexts [11, 10].
XI-E Edge Computing and Distributed Learning
The real-time and privacy-sensitive nature of crash detection necessitates a shift towards edge computing and distributed learning architectures. These approaches address the latency and data privacy challenges inherent in cloud-centric models.
XI-E1 Edge-Optimized Models
Deploying complex VLMs on resource-constrained edge devices (e.g., in-vehicle computers) requires significant model optimization. Techniques like pruning (removing redundant model weights), quantization (reducing the numerical precision of weights), and knowledge distillation (training a smaller model to mimic a larger one) are crucial for reducing model size and latency [69]. Research in edge-based video analytics for ITS has shown that such optimizations can enable real-time performance on embedded hardware [71]. For example, a quantized and pruned version of a model like MiniGPT-4 could be deployed on edge cameras for real-time crash detection, as its smaller footprint would meet the strict computational budget of in-vehicle systems.
XI-E2 Federated Learning for Privacy-Preserving Training
Federated learning offers a powerful paradigm for training models on decentralized data sources, such as data from a fleet of vehicles or a network of traffic cameras, without compromising privacy. In this approach, the model is trained locally on each device, and only the model updates (gradients), not the raw data, are sent to a central server for aggregation [72]. This is particularly valuable for crash detection, as it allows a global model to learn from a diverse range of driving scenarios and environments without requiring sensitive video data to be uploaded to the cloud, thus preserving the privacy of drivers.
XI-E3 Distributed Inference for Low Latency
For very large models, even after optimization, inference may be too slow for a single edge device. Distributed inference, where the computational load is split between the edge and the cloud, offers a solution. For example, an initial, low-latency analysis could be performed on the edge device to detect potential anomalies, and if a high-risk event is flagged, more detailed video data could be sent to a powerful cloud-based model for a more thorough analysis and descriptive summary [73]. This hybrid approach balances the need for low latency with the desire for deep, semantic understanding of crash events. A case study could involve deploying a lightweight variant of LAVAD on edge cameras to continuously monitor for anomalies, with its outputs triggering a more powerful VLM in the cloud for detailed causal analysis, with latency being a key evaluation metric on a dataset like CCD [37, 44].
XI-F Active Learning and Human-in-the-Loop (HITL)
To further address data scarcity and improve model reliability, active learning and Human-in-the-Loop (HITL) systems present a promising frontier. These approaches can create a virtuous cycle of data collection, annotation, and model refinement, making the entire system more efficient and trustworthy.
XI-F1 Active Learning for Efficient Annotation
Active learning can significantly reduce the manual annotation burden by intelligently selecting the most informative data for labeling [74]. For crash detection, an active learning system could prioritize video clips where the model is most uncertain about its predictions [75]. For example, after an initial model is trained on a small subset of DAD or CADP, it could be used to score a large, unlabeled pool of traffic videos. The videos with the highest uncertainty would then be sent to human annotators. This targeted annotation strategy ensures that labeling efforts are focused on the most challenging and valuable examples, accelerating model improvement and reducing the cost of creating large-scale datasets like DAD or CADP [38, 39].
XI-F2 Human-in-the-Loop for Validation and Refinement
HITL systems are essential for safety-critical applications, where the cost of an error is high. In the context of crash detection, a HITL system would involve human experts validating the outputs of models like VERA or CrashLLM [76]. This is particularly important for correcting ”hallucinations” or factual inaccuracies in LLM-generated explanations of crash events. By providing corrective feedback, human experts can iteratively refine the model’s performance and ensure its outputs are reliable and trustworthy. This process also generates a valuable dataset of human-verified crash descriptions, which can be used for further fine-tuning. Evaluation metrics for such a system could go beyond standard captioning scores like BLEU to include human agreement scores and measures of annotation efficiency, providing a more holistic assessment of the system’s performance.
The proposed future directions—synthetic training data, video-grounded QA benchmarks, fine-tuned VLMs, and multilingual and low-resource models—address the core challenges of LLM-based crash detection systems. Synthetic data generation can mitigate data scarcity, building on datasets like BDD100K and CADP [40, 39]. QA benchmarks will enhance explainability, complementing methods like VERA [29]. Fine-tuned VLMs will improve multimodal alignment, leveraging architectures like TimeSformer [52]. Multilingual and low-resource models will ensure global applicability, aligning with real-time needs [37]. These directions, supported by foundational work in vision-language pre-training [23, 24, 28] and temporal modeling [51, 49], pave the way for robust, interpretable, and accessible crash detection systems, advancing the field beyond the methods discussed in Section VI [11].
XII Conclusion
The advent of Large Language Models (LLMs) and Vision-Language Models (VLMs) has transformed crash detection from video data, enabling high-level reasoning that bridges visual observations with linguistic descriptions. This survey comprehensively reviewed advancements in LLM-based crash detection from 2023 to 2024, covering models, datasets, taxonomies, challenges, and future directions, with the goal of guiding researchers toward building reliable and explainable multimodal systems for safety-critical applications in intelligent transportation systems [11, 33]. By synthesizing recent developments, this work underscores the potential of LLMs to revolutionize traffic safety while identifying critical areas for further innovation.
Section VI analyzed state-of-the-art methods, such as CrashLLM, VERA, Holmes-VAD, and Video-LLaMA, which leverage multimodal fusion strategies (e.g., early, late, cross-attention) to achieve high accuracy and explainability [33, 29, 56, 27]. These methods outperformed classical approaches, such as CNN-based systems, by integrating visual and textual data to address complex scenarios involving occlusions and variable lighting [3, 6]. For instance, CrashLLM’s generative capabilities produce detailed crash descriptions, achieving over 92% accuracy, while VERA’s verbalized learning enhances post-event analysis with 86.55% AUC [33, 29]. The survey also introduced a taxonomy of fusion strategies and LLM roles (e.g., passive, generative), providing a structured framework for understanding these advancements [11, 27].
Key datasets, including BDD100K, CADP, and UCF-Crime, have been instrumental in training these models, though their limitations—such as limited crash-specific videos and coarse annotations—pose challenges [40, 39, 41]. The recent emergence of large-scale, finely annotated datasets like SHRP 2 NDS [42] is a critical step forward. Section VII detailed these obstacles, highlighting data scarcity, multimodal alignment difficulties, reasoning errors (e.g., hallucination), and real-time computational constraints [38, 30, 37]. For example, the Dashcam Accident Dataset (DAD) contains only 1,500 videos, insufficient for large-scale VLM training, while real-time systems like Video-LLaMA struggle with latency requirements of 30–100 ms [38, 27]. These challenges underscore the need for innovative solutions to enhance robustness and practicality.
Section XI proposed four directions to address these issues: generating synthetic training data using simulators like CARLA, developing video-grounded QA benchmarks, fine-tuning VLMs for crash-specific scenarios, and creating multilingual and low-resource models [40, 32, 45]. These strategies aim to mitigate data scarcity, improve temporal alignment, enhance explainability, and ensure global accessibility. For instance, synthetic data can augment datasets like CADP, while QA benchmarks can enable interactive crash analysis [39]. Fine-tuning with techniques like LoRA and multilingual pre-training with models like MiniGPT-4 can further advance performance and equity [32, 45].
The implications of these advancements are profound for safety-critical applications, such as autonomous driving, traffic surveillance, and accident response. Methods like Holmes-VAD and LAVAD demonstrate the potential for unbiased and training-free crash detection, while TrafficVLM highlights the power of dense captioning for post-event analysis [56, 37, 43]. Foundational models, including CLIP, Flamingo, and GPT-4, have laid the groundwork for robust feature extraction and reasoning, supported by evaluation metrics like BLEU, METEOR, and CIDEr [23, 26, 30, 53, 54, 55]. Temporal modeling architectures, such as TimeSformer, SlowFast, and ConvLSTM, further enhance video understanding, addressing dynamic crash scenarios [52, 46, 51].
This survey encourages researchers to build on these foundations, leveraging insights from recent methods and proposed directions to develop multimodal systems that are accurate, interpretable, and deployable in real-world settings. By addressing challenges through innovations like synthetic data and fine-tuned VLMs, the field can move toward standardized benchmarks and efficient architectures, as seen in works like CLIP4Clip and scaling approaches [28, 24]. Ultimately, we hope this survey serves as a catalyst for advancing LLM-based crash detection, fostering safer and more intelligent transportation systems [11, 10].

### v2

@@ VII-A Data Scarcity
Current crash datasets (DAD: 1,500 videos) are insufficient for large-scale multimodal models due to crash rarity and annotation complexity [61, 38]. Solutions include synthetic data generation via CARLA/SUMO simulators, self-supervised pre-training on unlabeled driving data, and federated learning for privacy-preserving collaboration.
@@ VII-B Multimodal Alignment
Temporal synchronization between video and language is challenging under occlusions and variable framerates [54]. Cross-attention mechanisms address this but increase computational complexity [55].
@@ VII-C Reasoning and Explainability
LLM hallucination risks factually incorrect crash reports in safety-critical applications [17]. Robust grounding techniques and causal intervention modules are needed to ensure visual-textual consistency [54, 51].
@@ VII-D Real-time Constraints
LLM-based models for video analysis often face inference latencies in the range of hundreds of milliseconds to seconds, which can exceed the 30–100 ms per-frame budget typically required for real-time safety-critical applications [73, 88]. While exact timings are rarely reported in crash-specific studies, the computational overhead of 7B+ parameter models is widely acknowledged. Potential solutions discussed in the broader literature include model quantization for reduced compute cost, hierarchical pipelines where lightweight detectors filter candidate events before LLM-based reasoning, and edge–cloud hybrid architectures for distributing workloads [17, 87].
@@ VII-E Ethical and Fairness Considerations
Dataset biases and privacy concerns require stratified data construction, bias auditing, differential privacy, and clear accountability frameworks with explainable AI for safety-critical applications.
VII-F Robustness and Failure Analysis
Models show significant performance drops (12-16%) in out-of-distribution scenarios [89]. Critical Failure Modes: (1) Occlusion failures where 40% of crashes are missed when key vehicles are partially obscured, (2) Weather degradation with 25% accuracy drop in rain/snow conditions, (3) Lighting failures in dawn/dusk scenarios reducing detection by 30%, (4) Adversarial vulnerability where imperceptible perturbations can cause 60% false negatives, raising security concerns for malicious attacks. Solutions include domain generalization, adversarial training, OOD detection mechanisms, and systematic cross-dataset evaluation [90, 91].
VII-F1 Augmentation with Synthetic Data
@@
VII-G Domain Adaptation and Generalization Strategies
Current models show significant performance drops (12-16% on average) when deployed in new environments. We propose several strategies to address this critical limitation:
VII-G1 Geographic Adaptation
Multi-region training includes data from diverse geographic regions (urban/rural, different countries) in training sets. Style transfer uses domain adaptation techniques to transfer models trained on one region to another. Local fine-tuning develops protocols for rapid adaptation using 100-500 local examples.
VII-G2 Weather and Lighting Robustness
Adversarial weather training augments training data with synthetic weather conditions (rain, snow, fog, night). Multi-spectral integration combines visible and infrared cameras for all-weather operation. Temporal consistency uses video sequences rather than single frames to maintain performance in poor visibility.
VII-G3 Camera and Hardware Adaptation
Resolution invariance trains models to handle varying camera resolutions and viewing angles. Calibration-free methods develop approaches that don’t require precise camera calibration. Hardware-agnostic deployment ensures models work across different edge computing platforms.
These challenges reveal a significant deployment gap between laboratory results and real-world requirements, highlighting the fundamental mismatch between data-hungry LLMs and resource-constrained transportation systems [62, 60].
VIII Integration with Autonomous Driving Ecosystems
Figure 8: Integration of an LLM-based crash detection module within a standard Autonomous Vehicle (AV) perception-planning-action pipeline. The model provides semantic risk assessment to the planning module, enabling more nuanced and proactive vehicle control.
@@ VIII Integration with Autonomous Driving Ecosystems
VIII-A AV Pipeline Integration
A standard AV software architecture follows a perception-planning-action pipeline [93]. LLM-based crash detection serves as an advanced component of the perception stack. The outputs from a model like VERA or CrashLLM—such as a textual description (imminent rear-end collision due to sudden braking”) or a classification of risk—can be fed directly into the AV’s decision-making or planning module. This allows the AV to move beyond simple object detection and react to a nuanced understanding of the scene’s dynamics. For example, a high-risk prediction could trigger an emergency braking maneuver or a swerving action, demonstrating a direct link between semantic crash understanding and vehicle control.
VIII-B Multi-Sensor Fusion
While cameras are the primary input for the VLM-based methods discussed, their fusion with other sensor modalities like LiDAR and radar is crucial for building a robust perception system for AVs. LiDAR provides precise 3D geometric data, which is less affected by adverse weather, while radar offers reliable velocity information. A multi-sensor fusion approach would combine the semantic, contextual understanding from an LLM/VLM with the geometric and velocity data from LiDAR/radar [94]. For example, a VLM might identify a “pedestrian looking at their phone,” and this semantic flag could be fused with LiDAR data confirming their trajectory towards the road, leading to a more reliable and timely pre-emptive braking action than either sensor could achieve alone. Extending frameworks like Video-LLaMA to incorporate these additional data streams is a promising research direction.
VIII-C Real-time Performance and Latency
@@
VIII-D Crash Anticipation as a Proactive Safety Measure
@@
IX Future Directions
Based on the challenges and advancements discussed, this section proposes key future directions to enhance robustness, scalability, and applicability of LLM-based crash detection systems.
IX-A Synthetic Training Data
Data scarcity limits LLM generalization [61, 62]. Synthetic training data from CARLA/SUMO simulators can generate diverse crash scenarios with textual narratives, augmenting datasets like DAD and UCF-Crime [63, 64]. Hybrid synthetic-real datasets with domain adaptation techniques offer scalable solutions for rare crash scenarios [53, 51].
IX-B Video-grounded QA Benchmarks
@@
IX-C Fine-tuned VLMs
@@
While pre-trained VLMs like BLIP-2 and Flamingo offer strong generalization, their performance in crash-specific scenarios is limited by the lack of domain-specific fine-tuning [19, 18]. Fine-tuning VLMs on crash-specific scenes and annotations can enhance their ability to detect and describe complex crash events, addressing multimodal alignment issues [54]. For instance, fine-tuning Video-LLaMA on datasets like DAD or BDD100K, with annotations for crash types, severity, and temporal dynamics, can improve temporal synchronization and feature extraction [38, 63]. Techniques like LoRA (Low-Rank Adaptation) enable efficient fine-tuning of large models, reducing computational overhead while adapting to crash-specific contexts [52]. Recent work on VERA demonstrates the benefits of fine-tuning LLaVA-1.5 for explainable crash detection, achieving high AUC and BLEU scores [54]. Future efforts should focus on curating crash-specific annotations, incorporating temporal models like TimeSformer or SlowFast networks to capture dynamic events.
X Conclusion
This survey examined LLM/VLM-based video crash detection and highlighted the persistent gap between promising research results and real-world deployment constraints. Recent systems show strong capabilities in multimodal understanding, but many works still underreport latency, struggle with cross-domain robustness, and rely on datasets that are relatively small or weakly annotated for safety-critical use. See, e.g., LLM/VLM methods such as Video-LLaMA for video understanding, and training-free approaches like LAVAD for anomaly detection [38, 60].
X-A Key Findings
  1. 1.
    Deployment Gap: High-capacity VLMs can achieve strong accuracy on curated benchmarks, but real-time, edge-friendly latency is rarely reported and remains a major barrier for on-device ITS deployment. Training-free pipelines (e.g., LAVAD) reduce data/engineering overhead but depend on captioning backends whose speed varies across hardware and models [38, 60].
  2. 2.
    Generalization Challenge: Across the broader video anomaly literature, models often degrade when moved off their source domain/dataset; works frequently note domain specificity and the need for adaptation when transferring between datasets such as UCF-Crime, CADP, and DAD [10, 62, 61].
  3. 3.
    Data Bottlenecks: Several widely used datasets for crash/anomaly detection remain modest in size. For example, DAD contains $\sim$620 dashcam videos (with positive/negative splits), and UCF-Crime has $\sim$1,900 untrimmed surveillance videos across 13 anomaly classes [61, 10].
X-B Critical Recommendations
For Researchers: Design efficiency-first architectures and report latency alongside accuracy; incorporate domain adaptation/evaluation across datasets; and consider training-free or weakly-supervised regimes where appropriate. Privacy-preserving and federated paradigms are increasingly relevant for multi-site video data [96].
For Practitioners: Start with lightweight or training-free baselines (e.g., LAVAD-style pipelines) before committing to heavy retraining; adopt hierarchical processing (quick filters $\to$ deeper reasoning on suspected segments); plan for local adaptation and monitoring; and keep human oversight in the loop for safety-critical decisions [60].
X-C Synthesis
We observe: (i) Reasoning vs. Deployability: richer temporal/contextual reasoning often coincides with higher compute cost; (ii) Source–Target Drift: models tuned to one dataset typically require adaptation to new camera placements, traffic patterns, and artifact distributions; (iii) System Design Tiers: in practice, teams converge on edge-first filters with selective offloading to stronger backends when latency budgets permit. These are qualitative patterns—reporting and standardized latency benchmarks remain an open need [38, 96].
X-D The Path Forward
Progress hinges on (a) stronger temporal modeling under efficiency constraints (e.g., video transformers like TimeSformer/ViViT), (b) robust multilingual and low-resource operation via VLM alignment/fine-tuning, (c) privacy-aware training via federated or distributed learning where data cannot leave premises, and (d) human-in-the-loop active learning for targeted data curation and assurance [38, 96].